OpenStudy (anonymous):

Evaluate \[\int\limits_{-1}^{0}\sqrt{1-x ^{2}}dx \] by interpreting it in terms of area

OpenStudy (eyust707):

i would draw a triangle

OpenStudy (anonymous):

it's the top half a circle with diameter 1

OpenStudy (eyust707):

theres your answer =)

OpenStudy (eyust707):

actually its r= 1

OpenStudy (eyust707):

and x goes from -1 to 0 so its a quater circle

OpenStudy (jamesj):

If \( y = \sqrt{1-x^2} \), then \( y^2 = 1 -x^2 \) or \( x^2 + y^2 = 1 \) Hence it is indeed a circle of radius r = 1 and it's the quarter lying in the 2nd quadrant.

OpenStudy (eyust707):